What monthly payment is required to amortize a loan of $35,000 over 10 yr if interest at the rate of 15%/year is charged on the unpaid balance and interest calculations are made at the end of each month? (Round your answer to the nearest cent.)
Answer #1
i = .15/12 = .0125
n = 10x12 = 120

35000 = pay(1 - 1.0125^-120)/.0125
35000 = pay(61.98284725
payment = $564.67
answered by: soliman